# An Asymptotically Optimal Coordinate Descent Algorithm for Learning Bayesian Networks from Gaussian Models

### v1

@@ 4.1. Convergence and Optimality Guarantees
Here, the first equality follows from appealing to (8), and noting that $\hat{\Gamma}_{vu}=0$ and that $(\hat{\Sigma}\hat{\Gamma})_{uv}=0$ according to item ii. of Lemma 8; the second equality follows from item iii. of Lemma 8; the final equality follows from noting that $\hat{\Gamma}_{ui}=0$ for $i\in F_{vu}$.
For each $i\in F_{vu}$, Figure 1 (left) represents the relationships between the nodes $u,v,i$. Here, the directed edge from $u$ to $v$ from the constraint $(u,v)\in\hat{E}$ is represented by a split line, the directed edge from $v$ to $i$ from the constraint $i\in F_{vu}$ is represented by a solid line, and the directed edge that is disallowed due to the constraint $i\in F_{vu}$ is represented via a cross-out solid line.
Since there is a directed path from $u$ to $i$, to avoid a cycle, a directed path from $i$ to $u$ cannot exist. Thus, adding the edge from $u$ to $i$ to $\hat{E}$ does not violate acyclicity and the fact that it is missing is due to $\lambda^{2}>A_{ui}^{2}/(4\hat{\Sigma}_{uu})$ according to Proposition 1. Then, appealing to Lemma 10, we conclude that with probability greater than $1-\mathcal{O}(1/n)$: $|(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{vu}|\leq\sum_{i\in{F}_{vu}}\frac{1}{2}|\hat{\Gamma}_{vi}|2\lambda(\hat{\Sigma}_{u,u})^{1/2}\leq 4\lambda\bar{\kappa}d_{\text{max}}$. In other words, in this setting, $|(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{vu}|\to 0$ as $n\to\infty$.
@@
$(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{uv}$ $=\hat{\Gamma}_{uu}\left(\frac{1}{2}A_{vu}+\hat{\Gamma}_{vu}\hat{\Sigma}_{vv}\right)+\sum_{i\in F_{uv}}\hat{\Gamma}_{ui}\left(\frac{1}{2}A_{vi}+\hat{\Gamma}_{vi}\hat{\Sigma}_{vv}\right)$
(9) $=\sum_{\begin{subarray}{c}i\in F_{uv}\\ \cup\{u\}\end{subarray}}\frac{\hat{\Gamma}_{ui}A_{vi}}{2}.$
@@
For each $i\in F_{uv}$, Figure 1 (middle) represents the relations between nodes $u,v$, and $i$. Here, due to the condition of Setting II, nodes $u$ and $v$ are not connected by an edge, so this is displayed by a solid crossed-out undirected edge. Furthermore, the directed edge from $u$ to $i$ from the constraint $i\in F_{uv}$ is represented via a solid directed edge, the directed edge $v$ to $i$ that is disallowed due to the constraint $i\in F_{uv}$ is represented via a cross-out solid line. Finally, the directed edge $u$ to $v$ that is disallowed due to acyclicity is represented via a crossed-out dashed line.
Since adding the directed edge $(u,v)$ to $\hat{E}$ creates a cycle, then we have the following implications: i. adding $(v,u)$ to $\hat{E}$ does not violate acyclicity (as both edges $u\to v$ and $v\to u$ cannot simultaneously create cycles) and ii. there must be a directed path from $v$ to $u$. Implication i. allows us to conclude that $\hat{\Gamma}_{vu}$ must be equal to zero due to the condition $4\hat{\Sigma}_{vv}\lambda^{2}>A_{vu}^{2}$ from Proposition 1. Combining implication ii. and the fact that there is a directed edge from $u$ to $i$ in $\hat{E}$ allows us to conclude that there cannot be a directed path from $i$ to $v$ as we would be creating a direct path from $u$ to itself. Thus, the fact that the directed edge $(v,i)$ is not in $\hat{E}$, or equivalently that $\hat{\Gamma}_{vi}=0$, is due to $4\hat{\Sigma}_{vv}\lambda^{2}>A_{vi}^{2}$ according to Proposition 1. From (9) and Lemma 10, we conclude with probability greater than $1-\mathcal{O}(1/n)$, $|(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{uv}|\leq 4\bar{\kappa}\lambda(1+d_{\text{max}})$. In other words, in this setting, $|(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{uv}|\to 0$ as $n\to\infty$.
@@
For each $i\in F_{uv}$, Figure 1 (right) represents the relations between the nodes $u,v$, and $i$ Here, due to the condition of Setting II, nodes $u$ and $v$ are not connected by an edge, so this is displayed by a solid crossed-out undirected edge. Furthermore, the directed edge from $u$ to $i$ from the constraint $i\in F_{uv}$ is represented via a solid directed edge, the directed edge $v$ to $i$ that is disallowed due to the constraint $i\in F_{uv}$ is represented via a cross-out solid line.
@@
Putting Settings I–II together, we have shown that as $n\to\infty$, the nonzero entries in the off-diagonal of $\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}$ specify a DAG. Furthermore, according to item i. of Lemma 8, the diagonal entries of this matrix are equal to one. As stated earlier, this then allows us to conclude that $-\log\det(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})\to 0$ as $n\to\infty$, and consequently that $f(\hat{\Gamma})-f(\hat{\Gamma}^{\mathrm{opt}})\to 0$.
uvi×$\times$
uvi×$\times$$\times$$\times$
uvi×$\times$$\times$
Figure 1. Left: scenario for Setting I, middle: scenario for setting II.1, and right: scenario for setting II.2; solid directed edges represent directed edges that are assumed to be in the estimate $\hat{E}$, crossed out solid directed edges represent directed edges that are assumed to be excluded in the estimate $\hat{E}$, crossed out solid undirected edges indicate that the corresponding nodes are not connected in $\hat{E}$, and crossed out dotted directed edge indicates that the edge is not present in $\hat{E}$ as adding it would create a cycle.
Part 2) Using the proof of Theorem 7 part i), we can immediately conclude that the matrix $\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}_{n}$ can be decomposed as the sum $N+\Delta$. Here, the off-diagonal entries of $N$ specify a DAG, with ones on the diagonal and under the assumption on $n$, with probability greater than $1-\mathcal{O}(1/n)$, $\|\Delta\|_{\infty}\leq 4\bar{\kappa}(1+d_{\text{max}})\lambda$ with zeros on the diagonal of $\Delta$. Consequently, $\|\Delta\|_{\infty}\leq 4m\bar{\kappa}^{2}(1+d_{\text{max}})\lambda$. Furthermore, by Lemma 10, we get $\sigma_{\text{min}}(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}_{n})\geq\sigma_{\text{min}}(\hat{\Gamma})^{2}\sigma_{\text{min}}(\hat{\Sigma})\geq\underline{\kappa}^{4}/4$. The reverse triangle inequality yields $\sigma_{\text{min}}(N)\geq\underline{\kappa}^{4}/4-4m\bar{\kappa}^{2}(1+d_{\text{max}})\lambda$. Consider any matrix $\bar{N}$ with $|\bar{N}_{ij}-N_{ij}|\leq|\Delta_{ij}|$. Using the reverse triangle inequality again, we get $\sigma_{\text{min}}(\bar{N})\geq\underline{\kappa}^{4}-8m\bar{\kappa}^{2}(1+d_{\text{max}})\lambda$ with probability greater than $1-\mathcal{O}(1/n)$. By the assumption on the sample size, $\bar{N}$ is invertible, and so we can use first-order Taylor series expansion to obtain $-\log\det(N+\Delta)=-\log\det(N)-\mathrm{tr}(\bar{N}^{-1}\Delta)$. Since $\log\det(N)=0$, we obtain the bound $-\log\det(N+\Delta)\leq-\mathrm{tr}(\bar{N}^{-1}\Delta)\leq\|\bar{N}^{-1}\|_{2}\|\Delta\|_{\star}$ with $\|\cdot\|_{\star}$ denoting the nuclear norm. Thus, $-\log\det(N+\Delta)\leq\|\bar{N}^{-1}\|_{2}\|\Delta\|_{\star}\leq\frac{m}{\sigma_{\text{min}}(\bar{N})}\|\Delta\|_{2}\leq\frac{4m^{2}\bar{\kappa}^{2}(1+d_{\text{max}})\lambda}{\underline{\kappa}^{4}/4-8m\bar{\kappa}^{2}(1+d_{\text{max}})\lambda}$. As $\lambda=\mathcal{O}(\log{m}/n)$, by the assumption on the sample size, $f(\hat{\Gamma})-f(\hat{\Gamma}^{\mathrm{opt}})\leq\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log{m}/n})$. ∎
4.2. Statistical Consistency Guarantees
Recall from Section 2.1 that there is typically multiple SEMs that are compatible with the distributions $\mathcal{P}^{\star}$. Each equivalent SEM is specified by a DAG; this DAG defines a total ordering among the variables. Associated to each ordering $\pi$ is a unique structural equation model that is compatible with the distribution $\mathcal{P}^{\star}$. We denote the set of parameters of this model as $(\tilde{B}^{\star}(\pi),\tilde{\Omega}^{\star}(\pi))$. For the tuple $(\tilde{B}^{\star}(\pi),\tilde{\Omega}^{\star}(\pi))$, we define $\tilde{\Gamma}^{\star}(\pi):=(I-\tilde{B}^{\star}(\pi)){\tilde{\Omega}}^{\star}(\pi)^{-1/2}$. We let $\Pi=\{\text{ordering }\pi:\mathrm{support}(\tilde{B}^{\star}(\pi))\subseteq E_{\text{super}}\}$. Throughout, we will use the notation $s^{\star}=\|B^{\star}\|_{\ell_{0}}$ and $\tilde{s}:=\tilde{s}^{\star}(\pi)=\|\tilde{B}^{\star}(\pi)\|_{\ell_{0}}$.
Assumption 3.
(Sparsity of every equivalent causal model) There exists some constant $\tilde{\alpha}$ such that for any $\pi\in\Pi$, $\|\tilde{B}^{\star}_{:j}(\pi)\|_{\ell_{0}}\leq\tilde{\alpha}\sqrt{n}/\log{m}$.
Assumption 4.
(Beta-min condition) There exist constants $10\leq\eta_{1}<$ 1 and $0<\eta_{0}^{2}<1-\eta_{1}$, such that for any $\pi\in\Pi$, the matrix $\tilde{B}^{\star}(\pi)$ has at least $(1-\eta_{1})\|\tilde{B}^{\star}(\pi)\|_{\ell_{0}}$ coordinates $k\neq j$ with $|\tilde{B}^{\star}_{kj}(\pi)|>\sqrt{\log{m}/n}(\sqrt{m/{s^{\star}}}\lor 1)/\eta_{0}$.
Assumption 5.
(Sufficiently large noise variances) For every $\pi\in\Pi$, $\mathcal{O}(1)\geq\min_{j}\allowbreak[\tilde{\Omega}^{\star}(\pi)]_{jj}\geq\mathcal{O}(\sqrt{s^{\star}\log{m}/n})$.
Assumption 6.
(Sufficiently sparse $B^{\star}$ and super-structure $E_{\text{super}}$) For every $i=1,2,\dots,m$, $\|B^{\star}_{:,i}\|_{\ell_{0}}\leq\alpha{n}/\log(m)$ and |$|\{j,(j,i)\in E_{\text{super}}\}|\leq\alpha{n}/\log(m)$.
Here, Assumptions 3-4 are similar to those in [23]. Assumption 5 is used to characterize the behavior of the early stopped estimate and is thus new relative to [23]. Assumption 6 ensures that the number of parents for every node both in the true DAG and the super-structure is not too large.
Next, we present our theorem on the finite-sample consistency guarantees of the coordinate descent algorithm. Throughout, we assume that we have obtained a solution after the algorithm has converged. We let $\mathrm{GAP}$ denote the difference between the objective value of the coordinate descent output and the optimal objective value of (3). We let $\hat{\Gamma}$ be a minimizer of (3).
Theorem 11.
Let $\hat{\Gamma},\hat{\Gamma}^{\mathrm{opt}}$ be the solution of Algorithm 1 and the optimal solution of (3), respectively. Suppose Assumptions 2-6 are satisfied with constants $\alpha,\tilde{\alpha},\eta_{0}$ sufficiently small. Let $\alpha_{0}:=\min\{4/m,0.05\}$. Then, for $\lambda^{2}\asymp\log{m}/n$, if $n/\log(n)\geq\mathcal{O}(m^{2}\log{m})$, with probability greater than $1-2\alpha_{0}$, there exists a $\pi$ such that
  1. (1)
    $\|\hat{\Gamma}-\hat{\Gamma}^{\mathrm{opt}}\|_{F}^{2}\leq\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log{m}/n})$,
  2. (2)
    $\|\hat{\Gamma}-\tilde{\Gamma}^{\star}(\pi)\|_{F}^{2}=\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log{m}/n})$, and $\|\tilde{\Gamma}^{\star}(\pi)\|_{\ell_{0}}\allowbreak\asymp s^{\star}$.
The proof relies on the following results.
Proposition 12.
(Theorem 3.1 of [23]) Suppose Assumptions 26 hold with constants $\alpha,\tilde{\alpha},\eta_{0}$ sufficiently small. Let $\hat{\Gamma}^{\mathrm{opt}}$ be any optimum of (3) with the constraint that $\mathrm{support}(\Gamma)\subseteq E_{\text{super}}$. Let $\pi^{\mathrm{opt}}$ be the associated ordering of $\hat{\Gamma}^{\mathrm{opt}}$ and $(\hat{B}^{\mathrm{opt}},\hat{\Omega}^{\mathrm{opt}})$ be the associated connectivity and noise variance matrix satisfying $\hat{\Gamma}^{\mathrm{opt}}={(I-\hat{B}^{\mathrm{opt}})\hat{\mathcal{K}}^{\mathrm{opt}}}^{-1/2}$. Then, for $\alpha_{0}:=(4/m)\land 0.05$ and $\lambda^{2}\asymp{\log{m}}/{n}$, we have, with a probability greater than $1-\alpha_{0}$, $\|\hat{B}^{\mathrm{opt}}-\tilde{B}^{\star}(\pi)\|_{F}^{2}+\|\hat{\Omega}^{\mathrm{opt}}-\tilde{\Omega}^{\star}(\pi^{\mathrm{opt}})\|_{F}^{2}=\mathcal{O}(\lambda^{2}s^{\star})$, and $\|\tilde{B}^{\star}(\pi)\|_{\ell_{0}}\asymp s^{\star}$.
Corollary 13 (Corollary 6 of [25]).
With the setup in Proposition 12,
$\left\|\hat{\Gamma}^{\mathrm{opt}}-\tilde{\Gamma}^{\star}(\pi)\right\|_{F}^{2}\leq\frac{16\max\{1,\|\tilde{B}^{\star}(\pi)\|_{F}^{2},\|\tilde{\Omega}^{\star}(\pi)^{-1/2}\|_{F}^{2}\}\lambda^{2}s^{\star}}{\min\{1,\min_{j}(\tilde{\Omega}^{\star}(\pi)_{jj})^{3}\}}.$
Proof of Theorem 11.
The proof is similar to that of [25] and we provide a short description for completeness. For notational simplicity, we let $\Gamma^{\star}:=\tilde{\Gamma}^{\star}(\pi)$ where $\pi$ is the permutation satisfying Proposition 12 and $\tilde{\Gamma}^{\star}$ defined earlier. From Theorem 7, we have that $0\leq f(\hat{\Gamma})-f(\hat{\Gamma}^{\mathrm{opt}})\leq\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log{m}/n})$. Let $\mathrm{GAP}=\mathcal{O}(\sqrt{d_{\mathrm{max}}^{2}m^{4}\log{m}/n})$. For a matrix $\Gamma\in\mathbb{R}^{m\times m}$, let $\mathcal{\ell}(\Gamma):=\sum_{i=1}^{m}-2\log(\Gamma_{ii})+\mathrm{tr}(\Gamma\Gamma^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}_{n})$. Suppose that $\|\hat{\Gamma}\|_{\ell_{0}}\geq\|\hat{\Gamma}^{\mathrm{opt}}\|_{\ell_{0}}$. Then, $\ell(\hat{\Gamma})-\ell(\hat{\Gamma}^{\mathrm{opt}})\leq\mathrm{GAP}$. On the other hand, suppose $\|\hat{\Gamma}\|_{\ell_{0}}\leq\|\hat{\Gamma}^{\mathrm{opt}}\|_{\ell_{0}}$. Then, $\ell(\hat{\Gamma})-\ell(\hat{\Gamma}^{\mathrm{opt}})\leq\mathrm{GAP}+\lambda^{2}\|\hat{\Gamma}\|_{\ell_{0}}\leq 2\mathrm{GAP}$. So, we conclude the bound $\ell(\hat{\Gamma})-\ell(\hat{\Gamma}^{\mathrm{opt}})\leq 2\mathrm{GAP}$.
For notational simplicity, we will consider a vectorized objective. Let $T\subseteq\{1,\ldots,m^{2}\}$ be indices corresponding to diagonal elements of an $m\times m$ matrix being vectorized. With abuse of notation, let $\hat{\Gamma}$, $\hat{\Gamma}^{\mathrm{opt}}$, and $\Gamma^{\star}$ be the vectorized form of their corresponding matrices. Then, Taylor series expansion yields

### v3

@@ 4.2 Optimality guarantees
Here, the first equality follows from appealing to (10), and noting that $\hat{\Gamma}_{vu}=0$ and that $(\hat{\Sigma}\hat{\Gamma})_{uv}=0$ according to item ii. of Lemma 8; the second equality follows from item i. of Lemma 8; the final equality follows from noting that $\hat{\Gamma}_{ui}=0$ for $i\in F_{vu}$.
For each $i\in F_{vu}$, Figure 1 (left) represents the relationships between the nodes $u,v,i$. Here, the directed edge from $u$ to $v$ from the constraint $(u,v)\in\hat{E}$ is represented by a dashed line, the directed edge from $v$ to $i$ from the constraint $i\in F_{vu}$ is represented by a solid line, and the directed edge that is disallowed due to the constraint $i\in F_{vu}$ is represented via a crossed-out solid line.
Since there is a directed path from $u$ to $i$, to avoid a cycle, a directed path from $i$ to $u$ cannot exist. Thus, adding the edge from $u$ to $i$ to $\hat{E}$ does not violate acyclicity and the fact that it is missing is due to $\lambda^{2}>A_{ui}^{2}/(4\hat{\Sigma}_{uu})$ according to Proposition 3. Then, appealing to Lemma 10, we conclude that with probability greater than $1-\mathcal{O}(1/n)$: $|(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{vu}|\leq\sum_{i\in F_{vu}}\frac{1}{2}|\hat{\Gamma}_{vi}|2\lambda(\hat{\Sigma}_{uu})^{1/2}\leq 4\lambda\bar{\kappa}^{2}d_{\text{max}}$. In other words, in this setting, $|(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{vu}|\to 0$ as $n\to\infty$.
@@
$(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{uv}$ $=\hat{\Gamma}_{uu}\left(\frac{1}{2}A_{vu}+\hat{\Gamma}_{vu}\hat{\Sigma}_{vv}\right)+\sum_{i\in F_{uv}}\hat{\Gamma}_{ui}\left(\frac{1}{2}A_{vi}+\hat{\Gamma}_{vi}\hat{\Sigma}_{vv}\right)=\sum_{\begin{subarray}{c}i\in F_{uv}\\ \cup\{u\}\end{subarray}}\frac{\hat{\Gamma}_{ui}A_{vi}}{2}.$ (11)
@@
For each $i\in F_{uv}$, Figure 1 (middle) represents the relations between nodes $u,v$, and ii. Here, due to the condition of Setting II, nodes $u$ and $v$ are not connected by an edge, which is displayed by a solid, crossed-out, undirected edge. Furthermore, the directed edge from $u$ to $i$ from the constraint $i\in F_{uv}$ is represented via a solid directed edge, the directed edge $v$ to $i$ that is disallowed due to the constraint $i\in F_{uv}$ is represented via a crossed-out solid line. Finally, the directed edge $u$ to $v$ that is disallowed due to acyclicity is represented via a crossed-out dashed line.
Since adding the directed edge $(u,v)$ to $\hat{E}$ creates a cycle, then we have the following implications: i. adding $(v,u)$ to $\hat{E}$ does not violate acyclicity (as both edges $u\to v$ and $v\to u$ cannot simultaneously create cycles) and ii. there must be a directed path from $v$ to $u$. Implication i. allows us to conclude that $\hat{\Gamma}_{vu}$ must be equal to zero due to the condition $4\hat{\Sigma}_{vv}\lambda^{2}>A_{vu}^{2}$ from Proposition 3. Combining implication ii. and the fact that there is a directed edge from $u$ to $i$ in $\hat{E}$ allows us to conclude that there cannot be a directed path from $i$ to $v$ as we would be creating a direct path from $u$ to itself. Thus, the fact that the directed edge $(v,i)$ is not in $\hat{E}$, or equivalently that $\hat{\Gamma}_{vi}=0$, is due to $4\hat{\Sigma}_{vv}\lambda^{2}>A_{vi}^{2}$ according to Proposition 3. From (11) and Lemma 10, we conclude with probability greater than $1-\mathcal{O}(1/n)$, $|(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{uv}|\leq\sum_{i\in F_{vu}\cup\{u\}}\frac{1}{2}|\hat{\Gamma}_{vi}|2\lambda(\hat{\Sigma}_{uu})^{1/2}\leq 4\bar{\kappa}\lambda(1+d_{\text{max}})$. In other words, in this setting, $|(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})_{uv}|\to 0$ as $n\to\infty$.
@@
For each $i\in F_{uv}$, Figure 1 (right) represents the relations between the nodes $u,v$, and ii. Here, due to the condition of Setting II, nodes $u$ and $v$ are not connected by an edge, so this is displayed by a solid crossed-out undirected edge. Furthermore, the directed edge from $u$ to $i$ from the constraint $i\in F_{uv}$ is represented via a solid directed edge, the directed edge $v$ to $i$ that is disallowed due to the constraint $i\in F_{uv}$ is represented via a crossed-out solid line.
@@
Putting Settings I–II together, we have shown that as $n\to\infty$, the nonzero entries in the off-diagonal of $\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma}$ specify a DAG. Furthermore, according to item iii. of Lemma 8, the diagonal entries of this matrix are equal to one. Since the off-diagonal pattern corresponds to a DAG, we can permute the rows and columns of this matrix to transform it into an upper triangular form without changing its determinant. Given that all diagonal entries are one, the determinant of the matrix is one, and hence $\log\det(\hat{\Gamma}\hat{\Gamma}^{\mathrm{\scriptscriptstyle T}}\hat{\Sigma})\to 0$ as $n\to\infty$, and, consequently, $f(\hat{\Gamma})-f(\hat{\Gamma}^{\mathrm{opt}})\to 0$.